1. Suppose the true proportion of voting Seattleites who support the $15 per hour wage increase is to be estimated. If we took a simple random sample of 805 likely voters in Seattle, and calculated the sample proportion who plan to vote in support of the proposition is 65%, then:a) what would be the standard error of the statistic?b) Calculate a 98% confidence interval for the proportion of voters in Seattle in support. Interpret.
1
Expert's answer
2019-11-01T09:51:52-0400

Let "\hat{p}" be the sample proportion.

When n is large and p is not close to zero or one, we can use the normal distribution to approximate the binomial.

The sampling distribution of "\hat{p}" is normal.

The mean of the sampling distribution is "p."

The standard deviation of the sampling distribution (called the standard error) is


"SE=\sqrt{{p(1-p) \over n}}"

Given that "p=0.65, n=805"

a)


"SE=\sqrt{{0.65(1-0.65) \over 805}}\approx0.016811"

b)

Confidence Level 98%: "z^*-" value"=2.33"

Confidence interval for "p" is


"\hat{p}\pm z^*\times SE"

"0.65\pm 2.33\times 0.016811\approx0.65\pm0.039"

"98\%\ CI=(0.611, 0.689)"